# Stochastic Population Update Provably Needs An Archive in Evolutionary Multi-objective Optimization

### v1

@@ 3.2 An Archive is Provably Helpful
In this section, we theoretically show that the limitation of SPU can be alleviated by using an archive. Once a new solution is generated, the solution will be tested if it can enter the archive. If there is no solution in the archive that dominates the new solution, then the solution will be placed in the archive, and meanwhile those solutions weakly dominated by the new solution will be deleted from the archive. Algorithmic steps incurred by adding an archive in SMS-EMOA are given as follows. In Algorithm 1, an empty set $A$ is initialized in line 1, the following lines are added after line 11:
@@
And the set $A$ instead of $P$ is returned in the last line.
We prove in Theorem 3 that if using an archive, a population size $\mu\geq 3$ is sufficient to guarantee the same running time bound of SMS-EMOA with SPU as Theorem 1. Note that Theorem 1 requires $\mu\geq(n-2k+4)/(1-p_{s})$, which implies that using an archive can allow a small population size and thus bring speedup. The proof idea is to divide the optimization procedure into three phases, where the first phase aims at finding the two boundary solutions of $F_{I}^{*}$ (i.e., a solution with $k$ 1-bits and a solution with $n-k$ 1-bits), the second phase aims at finding the two extreme Pareto optimal solutions $1^{n}$ and $0^{n}$, and the third phase aims at finding the remaining objective vectors on the Pareto front.
@@
We divide the running process into three phases. The first phase starts after initialization and finishes until the two boundary solutions of $F_{I}^{*}$ (i.e., a solution with $k$ 1-bits and a solution with $n-k$ 1-bits) are found; the second phase starts after the first phase and finishes when $1^{n}$ and $0^{n}$ are both found; the third phase starts after the second phase and finishes until the whole Pareto front is found.
For the first phase, we prove that the expected number of generations for finding a solution $\bm{x}$ with $n-k$ 1-bits is $O(\mu n\log n)$, and the same bound holds for finding a solution with $k$ 1-bits analogously. First, we show that the maximum $f_{1}$ value among the Pareto optimal solutions in $P\cup\{\bm{x}^{\prime\prime}\}$ will not decrease, where $\bm{x}^{\prime\prime}$ is the offspring solution generated in each iteration. Let $J$ denote the set of Pareto optimal solutions in $P\cup\{\bm{x}^{\prime\prime}\}$ with the maximum $f_{1}$ value. By definition, all solutions in $J$ are Pareto optimal, which means that they all belong to $R_{1}$ in the non-dominated sorting procedure. Furthermore, since all solutions in $J$ have the maximum $f_{1}$ value, there is at least one solution $\bm{x}^{*}\in J$ that is a boundary solution in $R_{1}$, which has the highest priority for SMS-EMOA to preserve. Since the survival probability $p_{s}$ satisfies $p_{s}\in[1/(\mu+1),(\mu-2)/(\mu+1)]$, the removed solution is selected from $\lfloor(\mu+1)\cdot(1-p_{s})\rfloor\geq 3$ solutions in $P\cup\{\bm{x}^{\prime\prime}\}$. Thus, if $\bm{x}^{*}$ is included in the survival selection, SMS-EMOA will directly preserve it, as it is one of the top two solutions. Note that for the bi-objective problem OneJumpZeroJump, there can be at most two boundary solutions in $R_{1}$. Thus, the maximum $f_{1}$ value among the Pareto optimal solutions in $P\cup\{\bm{x}^{\prime\prime}\}$ will not decrease.
During population initialization, as stated in Theorem 1, by the Chernoff bound and $n-2k=\Omega(n)$, there is at least one solution with the number of 1-bits between $k$ and $n-k$, with a probability of $1-\exp(-\Omega(n))^{\mu}$. Now, we consider the increase of the maximum $f_{1}$ value among the Pareto optimal solutions found. In each generation, SMS-EMOA selects the Pareto optimal solution $\bm{x}$ with the maximum $f_{1}$ value as a parent solution (whose probability is $1/\mu$), and generates a solution with more 1-bits if crossover is not performed (whose probability is $1-p_{c}$) and only one of the 0-bits in $\bm{x}$ is flipped by bit-wise mutation (whose probability is $((n-|\bm{x}|_{1})/n)\cdot(1-1/n)^{n-1}$). This implies that the probability of generating a solution with more than $|\bm{x}|_{1}$ 1-bits in one generation is at least
@@
Thus, the expected number of generations for increasing the maximum $f_{1}$ value to $n$, i.e., finding a solution with $n-k$ 1-bits, is at most $(1/(1-p_{c}))\cdot\sum_{i=k}^{n-k-1}e\mu n/(n-i)=O(\mu n\log n)$, where the equality holds by $p_{c}=\Theta(1)$. That is, the expected number of generations of the first phase is $O(\mu n\log n)$.
For the second phase, the analysis is the same as that of the second phase in Theorem 1. Before reaching the extreme Pareto optimal solution $1^{n}$, the boundary Pareto optimal solution with $n-k$ 1-bits is always the Pareto optimal solution with the maximum $f_{1}$ value and thus preserved in the population. Using SPU, the process of reaching $1^{n}$ can be constructed as jumping across the gap of the dominated solution set (i.e., $\{\bm{x}\mid|\bm{x}|_{1}\in[n-k+1..n-1]\}$) through $M$ stepping-stone solutions, and any failure during the intermediate jumps will result in restarting the process from the solution with $n-k$ 1-bits. By setting proper values of $M$, this process leads to an expected number of generations, $O(\mu n^{k}\cdot\min\{1,(e\ln C/k)^{k-1}\})$, where $C=e\mu/(p_{s}(1-p_{c}))$.
@@
Combining the three phases, the total expected number of generations is $O(\mu n\log n+\mu n^{k}\cdot\min\{1,(e\ln C/k)^{k-1}\}+\mu n\log n)=O(\mu n^{k}\cdot\min\{1,(e\ln C/k)^{k-1}\})$, where $C=e\mu/(p_{s}(1-p_{c}))$. The equality holds because $k\geq 2$ and $(en\ln C/k)^{k-1}\geq\log n$, as shown in the last paragraph in proof of Theorem 1. Thus, the theorem holds. ∎
@@
The smaller population size allows a larger probability of selecting inferior solutions that are close to Pareto optimal areas for reproduction, thus leading to speedup. When $k$ is limited, using only SPU may not bring acceleration. For example, when $k=\log n$, if using SPU with $p_{s}=1/2$, $\mu=2n$, and $p_{c}=1/2$, the expected running time of SMS-EMOA on OneJumpZeroJump is $O(\mu n^{k}\cdot\min\{1,(e\ln C/k)^{k-1}\})=O(n^{k+1})$ (where $C=e\mu/(p_{s}(1-p_{c}))=8en$), implying no acceleration compared to $O(\mu n^{k})=O(n^{k+1})$ without SPU Bian et al. (2024b). However, by adding an archive and reducing the population size $\mu$ to $5$, while keeping other settings unchanged, the bound reduces to $O(n^{k}\cdot(e\ln(20e)/k)^{k-1})$, implying a superpolynomial acceleration as $k=e\ln(8en)$. Note that we set $\mu=5$ to satisfy the condition of Theorem 3, i.e., to make $p_{s}=1/2\in[1/(\mu+1),(\mu-2)/(\mu+1)]$. On the other hand, when $k$ is large, using SPU alone can lead to exponential acceleration, and the addition of an archive can further enhance this acceleration. For example, when $k=n/8$, if using SPU with $p_{s}=1/2$, $\mu=2n$, and $p_{c}=1/2$, the expected running time is $O(\mu n^{k}\cdot(8e\ln(8en)/n)^{n/8-1})$, which implies an exponential acceleration compared to $O(\mu n^{k})$ without SPU. By adding an archive and setting $\mu$ to $5$, the bound reduces to $O(n^{k}\cdot(8e\ln(20e)/n)^{n/8-1})$, resulting in an improvement ratio of $\Theta(n(\ln(8en)/\ln(20e))^{n/8-1})$, i.e., an extra exponential acceleration.
@@ 4.1 NSGA-II with SPU
1:  $P\leftarrow\mu$ solutions uniformly and randomly selected from $\{0,\!1\}^{\!n}$ with replacement;
@@
7:  add $|Q|/2-|O|$ solutions in $R_{i}$ with the largest crowding distance into $O$;
@@
1:  $Q^{\prime}\leftarrow\lfloor|Q|\cdot(1-p_{s})\rfloor$ solutions uniformly and randomly selected from $Q$ without replacement;
@@
The expected running time of NSGA-II with SPU for solving the OneJumpZeroJump problem with $k>8e^{2}$ has been proved to be $O(\mu k(n/2)^{k})$ Bian et al. (2024b), which is better than that, i.e., $O(\mu n^{k})$, of the original NSGA-II Doerr and Qu (2023a). Here, we re-prove a tighter upper bound on the expected running time of NSGA-II with SPU for solving OneJumpZeroJump, as shown in Theorem 4. It is also more general, as it considers a survival probability $p_{s}\in[1/2\mu,1/2-1/o(\mu))$ and a crossover probability $p_{c}=O(1)$, rather than the specific probabilities $p_{s}=1/4$ and $p_{c}=0$ analyzed in Bian et al. (2024b).
@@
For NSGA-II solving OneJumpZeroJump with $n-2k=\Omega(n)$, if using SPU with survival probability $p_{s}\in[1/(2\mu),1/2-o(1/\mu))$, the crossover probability $p_{c}=O(1)$, and the population size $\mu\geq 4(n-2k+3)/(1-2p_{s})$, then the expected running time for finding the whole Pareto front is $O(\mu n^{k}\cdot\min\{1,(e\ln C/k)^{k-1}\})$, where $C=e/(p_{s}(1-p_{c}))$.
Before proving Theorem 4, we first present Lemma 5, which shows that given proper value of $\mu$, an objective vector on the Pareto front will always be maintained once it has been found.
@@
For NSGA-II solving OneJumpZeroJump, if using SPU with survival probability of $p_{s}\in[1/2\mu,1/2-1/o(\mu))$, and a population size $\mu$ such that $\mu\geq 4(n-2k+3)/(1-2p_{s})$, then an objective vector $\bm{f}^{*}$ on the Pareto front will always be maintained once it has been found.
@@
Suppose an objective vector $\bm{f}^{*}$ on the Pareto front is obtained by NSGA-II, i.e., there exists at least one solution in $Q$ (i.e., $P\cup P^{\prime}$ in line 14 of Algorithm 4) corresponding to the objective vector $\bm{f}^{*}$. Let $J$ denote the set of solutions in $P\cup P^{\prime}$ with objective vector $\bm{f}^{*}$, which are selected for competition in line 1 of Algorithm 6 (note that $J$ is a multi-set). Then, all solutions in $J$ belong to $R_{1}$ in the non-dominated sorting procedure, because these solutions are Pareto optimal. If the solutions in $J$ are sorted when computing the crowding distance in line 7 of Algorithm 6, the solution (denoted as $\bm{x}^{*}$) that is put in the first or the last position among solutions in $J$ will have a crowding distance greater than 00. For solutions with the same objective vector, they are sorted according to some objective function in the crowding distance assignment procedure, and only the first or the last one has crowding distance greater than 0. Since OneJumpZeroJump has two objectives, for each objective vector, there are at most $4$ solutions have crowding distance greater than 00. Considering that the Pareto front size of OneJumpZeroJump is $n-2k+3$, at most $4(n-2k+3)$ solutions in $R_{1}$ can have crowding distance greater than 00.
In Algorithm 6, $\lfloor|2\mu|\cdot(1-p_{s})\rfloor$ solutions are selected for competition, and $\lfloor\mu-2\mu\cdot p_{s}\rfloor\geq 4(n-2k+3)$ of them will not be removed. Note that $\bm{x}\in J$ is among the best $4(n-2k+3)$ solutions in $R_{1}$, implying $\bm{x}$ must be maintained in the next population. Thus, the lemma holds. ∎
@@
Next, we consider the second phase. We also use the $M+1$ jumps model of the analysis in Theorem 1. Specifically, by employing SPU, any solution (including dominated ones) can survive into the next generation with a probability of at least $p_{s}$. We also assume a “jump” to be an event where a solution $\bm{x}$ with $|\bm{x}|_{1}\in[n-k..n-1]$ is selected, and a new dominated solution closer to $1^{n}$ is generated and preserved. We refer to the dominated solutions along the multiple jumps as “stepping stones” and assume that $1^{n}$ can be reached through $M$ stepping stones. We consider $M+1$ consecutive jumps from $\bm{x}$ with $n-k$ 1-bits to the Pareto optimal solution $1^{n}$. Any failure during the intermediate jumps will result in restarting the process from the beginning. Note that in each generation, the probability of selecting a specific parent solution is changed from $1/\mu$ to $1$. Let $p_{i}$ ($11\leq i\leq M-$ 1) denotes the probability that, after selecting the $i$-th stepping-stone solution as a parent solution, the next stepping-stone solution is successfully generated and preserved. Let $p_{0}$ denotes the probability of selecting the solution $\bm{x}$ with $n-k$ 1-bits, generating the first stepping-stone solution, and preserving it. For the final jump, the Pareto optimal solution $1^{n}$ found must be preserved, and the jump occurs with a probability of $p_{M}$. Then, the probability of a successful trail of $M+1$ consecutive jumps is $\prod_{i=0}^{M}p_{i}$. Similar to the analysis in Eq. (4), we have
$\prod_{i=0}^{M}p_{i}\geq$ $\frac{p_{s}^{M}(1-p_{c})^{M+1}}{e^{M+1}n^{k}}\cdot\frac{k!}{k_{0}!\cdot k_{1}!\cdot k_{2}!\cdots k_{M}!}.$ (11)
@@
Combining the two phases, the total expected number of generations is $O(n\log n+n^{k}\cdot\min\{1,(e\ln C/k)^{k-1}\})=O(n^{k}\cdot\min\{1,(e\ln C/k)^{k-1}\})$, where $C=e/(p_{s}(1-p_{c}))$. The equality holds because $k\geq 2$ and $(en\ln C/k)^{k-1}\geq\log n$ as shown in the last paragraph in proof of Theorem 1. For NSGA-II, each generation requires $\mu$ fitness evaluations. Therefore, the expected number of fitness evaluations is $O(\mu n^{k}\cdot\min\{1,(e\ln C/k)^{k-1}\})$. Thus, the theorem holds. ∎
Bian et al. Bian et al. (2024b) analyzed NSGA-II solving OneJumpZeroJump with $k>8e^{2}$, $p_{s}=1/4$, and $p_{c}=0$, deriving an expected running time of $O(\mu k(n/2)^{k})$. In contrast, Theorem 4 gives a tighter bound $O(\mu n^{k}\cdot(e\ln(4e)/k)^{k-1})$, with an improvement ratio of $\Theta(k(k/(2e\ln(4e)))^{k-1})$, which can be exponential when $k$ is large, e.g., $k=n/8$. Our analysis leads to improvement because, in the second phase of finding the extreme Pareto optimal solution $1^{n}$, we consider an optimal number $M+1$ of jumps across the gap of the dominated solution set (i.e., $\{\bm{x}\mid|\bm{x}|_{1}\in[n-k+1..n-1]\}$), rather than a continuous sequence of $k$ jumps from $|\bm{x}|_{1}=n-k$ to $n$ in Bian et al. (2024b).

### v2

@@ 3.2 An Archive is Provably Helpful
In this section, we theoretically show that the limitation of SPU can be alleviated by using an archive. Once a new solution is generated, the solution will be tested if it can enter the archive. If there is no solution in the archive that dominates the new solution, then the solution will be placed in the archive, and meanwhile those solutions weakly dominated by the new solution will be deleted from the archive. Algorithmic steps incurred by adding an archive in SMS-EMOA are given as follows. In Algorithm 1, an empty set $A$ is initialized in line 1, and the following lines are added after line 11:
@@
The set $A$ instead of $P$ is returned in the last line.
We prove in Theorem 3 that if using an archive, a population size $\mu\geq 3$ is sufficient to guarantee the same running time bound of SMS-EMOA with SPU as Theorem 1. Note that Theorem 1 requires $\mu\geq(n-2k+4)/(1-p_{s})$, which implies that using an archive can allow a small population size and thus bring speedup.
@@
We divide the running process into three phases. The first phase starts after initialization and finishes until the two boundary solutions of $F_{I}^{*}$ (i.e., a solution with $k$ 1-bits and a solution with nkn-k 1-bits) are found; the second phase starts after the first phase and finishes when $1^{n}$ and $0^{n}$ are both found; the third phase starts after the second phase and finishes until the whole Pareto front is found.
For the first phase, we prove that the expected number of generations for finding a solution $\bm{x}$ with nkn-k 1-bits is $O(\mu n\log n)$, and the same bound holds for finding a solution with $k$ 1-bits analogously. First, we show that the maximum $f_{1}$ value among the Pareto optimal solutions in $P\cup\{\bm{x}^{\prime\prime}\}$ will not decrease, where $\bm{x}^{\prime\prime}$ is the offspring solution generated in each iteration. Let $J$ denote the set of Pareto optimal solutions in $P\cup\{\bm{x}^{\prime\prime}\}$ with the maximum $f_{1}$ value. By definition, all solutions in $J$ are Pareto optimal, which means that they all belong to $R_{1}$ in the non-dominated sorting procedure. Furthermore, since all solutions in $J$ have the maximum $f_{1}$ value, there is at least one solution $\bm{x}^{*}\in J$ that is a boundary solution in $R_{1}$, which has the highest priority for SMS-EMOA to preserve. Since the survival probability $p_{s}$ satisfies $p_{s}\in[1/(\mu+1),(\mu-2)/(\mu+1)]$, the removed solution is selected from $\lfloor(\mu+1)\cdot(1-p_{s})\rfloor\geq 3$ solutions in $P\cup\{\bm{x}^{\prime\prime}\}$. Thus, if $\bm{x}^{*}$ is included in the survival selection, SMS-EMOA will directly preserve it, as it is one of the top two solutions. Note that for the bi-objective problem OneJumpZeroJump, there can be at most two boundary solutions in $R_{1}$. Thus, the maximum $f_{1}$ value among the Pareto optimal solutions in $P\cup\{\bm{x}^{\prime\prime}\}$ will not decrease.
During population initialization, as stated in Theorem 1, by the Chernoff bound and $n-2k=\Omega(n)$, there is at least one solution with the number of 1-bits between $k$ and nkn-k, with a probability of $1-\exp(-\Omega(n))^{\mu}$. Now, we consider the increase of the maximum $f_{1}$ value among the Pareto optimal solutions found. In each generation, SMS-EMOA selects the Pareto optimal solution $\bm{x}$ with the maximum $f_{1}$ value as a parent solution (whose probability is $1/\mu$), and generates a solution with more 1-bits if crossover is not performed (whose probability is $1-p_{c}$) and only one of the 0-bits in $\bm{x}$ is flipped by bit-wise mutation (whose probability is $((n-|\bm{x}|_{1})/n)\cdot(1-1/n)^{n-1}$). This implies that the probability of generating a solution with more than $|\bm{x}|_{1}$ 1-bits in one generation is at least
@@
Thus, the expected number of generations for increasing the maximum $f_{1}$ value to $n$, i.e., finding a solution with nkn-k 1-bits, is at most $(1/(1-p_{c}))\cdot\sum_{i=k}^{n-k-1}e\mu n/(n-i)=O(\mu n\log n)$, where the equality holds by $p_{c}=\Theta(1)$. That is, the expected number of generations of the first phase is $O(\mu n\log n)$.
For the second phase, the analysis is the same as that of the second phase in Theorem 1. Before reaching the extreme Pareto optimal solution $1^{n}$, the boundary Pareto optimal solution with nkn-k 1-bits is always the Pareto optimal solution with the maximum $f_{1}$ value and thus preserved in the population. Using SPU, the process of reaching $1^{n}$ can be constructed as jumping across the gap of the dominated solution set (i.e., $\{\bm{x}\mid|\bm{x}|_{1}\in[n-k+1..n-1]\}$) through $M$ stepping-stone solutions, and any failure during the intermediate jumps will result in restarting the process from the solution with nkn-k 1-bits. By setting proper values of $M$, this process leads to an expected number of generations, $O(\mu n^{k}\cdot\min\{1,(e\ln C/k)^{k-1}\})$, where $C=e\mu/(p_{s}(1-p_{c}))$.
@@
Combining the three phases, the total expected number of generations is $O(\mu n\log n+\mu n^{k}\cdot\min\{1,(e\ln C/k)^{k-1}\}+\mu n\log n)=O(\mu n^{k}\cdot\min\{1,(e\ln C/k)^{k-1}\})$, where $C=e\mu/(p_{s}(1-p_{c}))$. The equality holds because $k\geq 2$ and $(en\ln C/k)^{k-1}\geq\log n$, as shown in the last paragraph in the proof of Theorem 1. Thus, the theorem holds. ∎
@@
The smaller population size allows a larger probability of selecting inferior solutions that are close to Pareto optimal areas for reproduction, thus leading to speedup. When $k$ is limited, using only SPU may not bring acceleration. For example, when $k=e\ln(8en)$, if using SPU with $p_{s}=1/2$, $\mu=2n$, and $p_{c}=1/2$, the expected running time of SMS-EMOA on OneJumpZeroJump is $O(\mu n^{k}\cdot\min\{1,(e\ln C/k)^{k-1}\})=O(n^{k+1})$ (where $C=e\mu/(p_{s}(1-p_{c}))=8en$), implying no acceleration compared to $O(\mu n^{k})=O(n^{k+1})$ without SPU Bian et al. (2024b). However, by adding an archive and reducing the population size $\mu$ to $5$, while keeping other settings unchanged, the bound reduces to $O(n^{k}\cdot(e\ln(20e)/k)^{k-1})$, implying a superpolynomial reduction in the upper bound as $k=e\ln(8en)$. Note that we set $\mu=5$ to satisfy the condition of Theorem 3, i.e., to make $p_{s}=1/2\in[1/(\mu+1),(\mu-2)/(\mu+1)]$. On the other hand, when $k$ is large, using SPU alone can lead to exponential acceleration, and the addition of an archive can further enhance this acceleration. For example, when $k=n/8$, if using SPU with $p_{s}=1/2$, $\mu=2n$, and $p_{c}=1/2$, the expected running time is $O(\mu n^{k}\cdot(8e\ln(8en)/n)^{n/8-1})$, which implies an exponential reduction in the upper bound compared to $O(\mu n^{k})$ without SPU. By adding an archive and setting $\mu$ to $5$, the bound reduces to $O(n^{k}\cdot(8e\ln(20e)/n)^{n/8-1})$, resulting in an improvement ratio of $\Theta(n(\ln(8en)/\ln(20e))^{n/8-1})$, i.e., an exponential reduction in the upper bound.
@@ 4.1 NSGA-II with SPU
[…]
@@
7: add |Q|/2|O||Q|/2-|O| solutions in $R_{i}$ with the largest crowding distance into $O$;
@@
[…]
@@
The expected running time of NSGA-II with SPU for solving the OneJumpZeroJump problem with $k>8e^{2}$ has been proven to be $O(\mu k(n/2)^{k})$ Bian et al. (2024b), which is better than that, i.e., $O(\mu n^{k})$, of the original NSGA-II Doerr and Qu (2023a). Here, we re-prove a tighter upper bound on the expected running time of NSGA-II with SPU for solving OneJumpZeroJump, as shown in Theorem 4. It is also more general, as it considers a survival probability $p_{s}\in[1/2\mu,1/2-o(1/\mu))$, rather than the specific probabilities $p_{s}=1/4$ analyzed in Bian et al. (2024b).
@@
For NSGA-II solving OneJumpZeroJump with $n-2k=\Omega(n)$, if using SPU with survival probability $p_{s}\in[1/(2\mu),1/2-o(1/\mu))$, the crossover probability $1-p_{c}=\Omega(1)$, and the population size $\mu\geq 4(n-2k+3)/(1-2p_{s})$, then the expected running time for finding the whole Pareto front is $O(\mu n^{k}\cdot\min\{1,(e\ln C/k)^{k-1}\})$, where $C=e/(p_{s}(1-p_{c}))$.
Before proving Theorem 4, we first present Lemma 5, which shows that given a proper value of $\mu$, an objective vector on the Pareto front will always be maintained once it has been found
@@
For NSGA-II solving OneJumpZeroJump, if using SPU with survival probability of $p_{s}\in[1/2\mu,1/2-o(1/\mu))$, and a population size $\mu$ such that $\mu\geq 4(n-2k+3)/(1-2p_{s})$, then an objective vector $\bm{f}^{*}$ on the Pareto front will always be maintained once it has been found.
@@
Suppose an objective vector $\bm{f}^{*}$ on the Pareto front is obtained by NSGA-II, i.e., there exists at least one solution in $Q$ (i.e., $P\cup P^{\prime}$ in line 14 of Algorithm 4) corresponding to the objective vector $\bm{f}^{*}$. Let $J$ denote the set of solutions in $P\cup P^{\prime}$ with objective vector $\bm{f}^{*}$, which are selected for competition in line 1 of Algorithm 6 (note that $J$ is a multi-set). Then, all solutions in $J$ belong to $R_{1}$ in the non-dominated sorting procedure, because these solutions are Pareto optimal. If the solutions in $J$ are sorted when computing the crowding distance in line 7 of Algorithm 6, the solution (denoted as $\bm{x}^{*}$) that is put in the first or the last position among solutions in $J$ will have a crowding distance greater than 0. For solutions with the same objective vector, they are sorted according to some objective function in the crowding distance assignment procedure, and only the first or the last one has crowding distance greater than 0. Since OneJumpZeroJump has two objectives, for each objective vector, there are at most $4$ solutions having crowding distance greater than 0. Considering that the Pareto front size of OneJumpZeroJump is n2k+3n-2k+3, at most 4(n2k+3)4(n-2k+3) solutions in $R_{1}$ can have crowding distance greater than 0.
In Algorithm 6, $\lfloor 2\mu\cdot(1-p_{s})\rfloor$ solutions are selected for competition, and $\lfloor\mu-2\mu\cdot p_{s}\rfloor\geq 4(n-2k+3)$ of them will not be removed. Note that $\bm{x}\in J$ is among the best 4(n2k+3)4(n-2k+3) solutions in $R_{1}$, implying $\bm{x}$ must be maintained in the next population. Thus, the lemma holds. ∎
@@
Next, we consider the second phase. We also use the analysis method of $M+1$ jumps in Theorem 1. Specifically, by employing SPU, any solution (including dominated ones) can survive into the next generation with a probability of at least $p_{s}$. We also assume a “jump” to be an event where a solution $\bm{x}$ with $|\bm{x}|_{1}\in[n-k..n-1]$ is selected, and a new dominated solution closer to $1^{n}$ is generated and preserved. We refer to the dominated solutions along the multiple jumps as “stepping stones” and assume that $1^{n}$ can be reached through $M$ stepping stones. We consider $M+1$ consecutive jumps from $\bm{x}$ with nkn-k 1-bits to the Pareto optimal solution $1^{n}$. Any failure during the intermediate jumps will result in restarting the process from the beginning. Note that in each generation, the probability of selecting a specific parent solution is changed from $1/\mu$ to $1$. Let $p_{i}$ ($1\leq i\leq M-1$) denote the probability that, after selecting the $i$-th stepping-stone solution as a parent solution, the next stepping-stone solution is successfully generated. Let $p_{0}$ denotes the probability of selecting the solution $\bm{x}$ with nkn-k 1-bits, and generating the first stepping-stone solution. For the final jump, the Pareto optimal solution $1^{n}$ found must be preserved, and the jump occurs with a probability of $p_{M}$. Then, the probability of a successful trail of $M+1$ consecutive jumps is $p_{s}^{M}\prod_{i=0}^{M}p_{i}$. Similar to the analysis in Eq. (4), we have
$p_{s}^{M}\prod_{i=0}^{M}p_{i}\geq$ $\frac{p_{s}^{M}(1-p_{c})^{M+1}}{e^{M+1}n^{k}}\cdot\frac{k!}{k_{0}!\cdot k_{1}!\cdot k_{2}!\cdots k_{M}!}.$ (11)
@@
Combining the two phases, the total expected number of generations is $O(n\log n+n^{k}\cdot\min\{1,(e\ln C/k)^{k-1}\})=O(n^{k}\cdot\min\{1,(e\ln C/k)^{k-1}\})$, where $C=e/(p_{s}(1-p_{c}))$. The equality holds because $k\geq 2$ and $(en\ln C/k)^{k-1}\geq\log n$ as shown in the last paragraph in the proof of Theorem 1. For NSGA-II, each generation requires $\mu$ fitness evaluations. Therefore, the expected number of fitness evaluations is $O(\mu n^{k}\cdot\min\{1,(e\ln C/k)^{k-1}\})$. Thus, the theorem holds. ∎
Bian et al. Bian et al. (2024b) analyzed NSGA-II solving OneJumpZeroJump with $k>8e^{2}$, $p_{s}=1/4$, and $p_{c}=0$, deriving an expected running time of $O(\mu\sqrt{k}(n/2)^{k})$. In contrast, Theorem 4 gives a tighter bound $O(\mu n^{k}\cdot(e\ln(4e)/k)^{k-1})$, with an improvement ratio of $\Theta(\sqrt{k}(k/(2e\ln(4e)))^{k-1})$, which can be exponential when $k$ is large, e.g., $k=n/8$. Our analysis leads to improvement because, in the second phase of finding the extreme Pareto optimal solution $1^{n}$, we consider an optimal number $M+1$ of jumps across the gap of the dominated solution set (i.e., $\{\bm{x}\mid|\bm{x}|_{1}\in[n-k+1..n-1]\}$), rather than a continuous sequence of $k$ jumps from $|\bm{x}|_{1}=n-k$ to $n$ in Bian et al. (2024b).